OpenStudy (anonymous):

-4(-3x) +5 (-7) which is equivalent to this expression?

OpenStudy (anonymous):

x. 5*x evaluates to 5x. 5*x+10 evaluates to 5x+10.

OpenStudy (anonymous):

....

OpenStudy (anonymous):

thx...

OpenStudy (anonymous):

I Will Give You Another One In A sec ok?!

OpenStudy (anonymous):

=-=

OpenStudy (anonymous):

Simplifying 2x + -10 = 10 + -3x Reorder the terms: -10 + 2x = 10 + -3x Solving -10 + 2x = 10 + -3x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '3x' to each side of the equation. -10 + 2x + 3x = 10 + -3x + 3x Combine like terms: 2x + 3x = 5x -10 + 5x = 10 + -3x + 3x Combine like terms: -3x + 3x = 0 -10 + 5x = 10 + 0 -10 + 5x = 10 Add '10' to each side of the equation. -10 + 10 + 5x = 10 + 10 Combine like terms: -10 + 10 = 0 0 + 5x = 10 + 10 5x = 10 + 10 Combine like terms: 10 + 10 = 20 5x = 20 Divide each side by '5'. x = 4 Simplifying x = 4

OpenStudy (anonymous):

PARTY!

OpenStudy (anonymous):

:D party! Thxxxx so much

OpenStudy (anonymous):

Wel!

OpenStudy (anonymous):

Distribut the 5 to the x and -2 and distribut the 2 to the 10 and x. You get 5x-10=20+2x…..subtract 5x from both sides to get -10=20-3x Subtract 20 from both sides to get -30=-3x……-30/-3=10......x=10 In Simplest For And When I Say This I Mean This One Is Easier To Understand!

OpenStudy (anonymous):

BRB!